A metal disk of radius 5.0 cm is mounted on a frictionless axle. Current can flow through the axle out along the disk, to a sliding contact on the rim of the disk. A forte at thegone iGetween the station ary electrical c onta ct and the rotating the surrent i5 3.0 A, the disk rotätes with constant angular velocity. Whats the frictiona 825 N This problem focuses on two concepts. One is the Faraday Law, the other is the conservation law of angular momentum. When the disk rotates at a uniform angular velocity, the net torque should be zero
Answer #1


magnetic moment u = I*A

torque due to magnetic field T1 = u*B*sintheta

theta = 90


T1 = I*A*B


torque due to frictional force T2 = -f*R

net torque = T1 + T2


but in constant angular motion net torque = 0

T1 + T2 = 0

I*A*B - f*R = 0

I*pi*R2*B - f*R = 0


f = I*pi*R*B

f = 3*pi*0.05*1.75


frictional force f = 0.825 N <<<--------ANSWER